CHANGE POINT ESTIMATION IN NON-MONOTONIC AGING MODELS

MURARI MITRA^{1} AND SUJIT K. BASU^{2}

1Division of Theoretical Statistics and Mathematics, Indian Statistical Institute,

203, B.T. Road, Calcutta 700 035, India 2Indian Institute of Management, Joka, Diamond Harbour Road, Post Office Box No. 16757, Calcutta 700 027, India

(Received March 14, 1994; revised December 5, 1994)

Abstract. The problem of estimating change points in various non-monoto-nic aging models is considered. A general methodology for consistent estimation of the change point is developed and applied to non-monotonic aging models based on the hazard rate function as well as on the mean residual life function.

Key words and phrases: Non-monotonic aging, NWBUE, IDMRL, BFR life distributions, change points.